OpenStudy (anonymous):

PLZZZZZZZZZZZ HELP! Melinda, Marsha and Joanna paid an extra $35 fee for a suitcase on an airline. Of its contents, 12 lbs was Melinda's, 17lbs was Marsha's, and 21lbs was Joanna's. How should they divide the fee.

OpenStudy (anonymous):

lets try that again \[12+17+21=50\] so each pound should cost \[\frac{35}{50}=.7\]

OpenStudy (anonymous):

melinda should pay \(12\times .7=\$8.40\) etc

OpenStudy (anonymous):

Oh ok. Thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

y, wht?